OpenStudy (anonymous):

If the y intercept remains the same and the slope is doubled, what will be the x intercept? This is with y=-5/2x+3.

OpenStudy (anonymous):

Double the slope. So -5/2 becomes -5. Then y=-5x+3 Set y=0 and solve for x.

OpenStudy (anonymous):

A sphere has a volume of 7235 cubic centimeters and a surface area of 1809 square centimeters. What is the radius of the sphere?

OpenStudy (anonymous):

Well you can use either the surface area or volume to solve for radius. V=(4/3)*pi*r^3 SA = 4*pi*r^2. You are given V and SA. r is then the only unknown.